OpenStudy (anonymous):

4^2x-3=32 solve exponential equation

OpenStudy (xapproachesinfinity):

is it like this \(\normalsize\color{blue}{ 4^{2x-3} = 32 }\)

OpenStudy (xapproachesinfinity):

So do you know how to use exponential and logarithm, like \(\normalsize\color{blue}{ log{2^x=x} }\)

OpenStudy (anonymous):

Yes @xapproachesinfinity

OpenStudy (xapproachesinfinity):

32 you can write as a power \(\normalsize\color{blue}{ 32= 2^5 }\)

OpenStudy (xapproachesinfinity):

I suggest that you rewrite \(\normalsize\color{blue}{ 4^{2x=3}=2^{2(2x-3)} }\)

OpenStudy (xapproachesinfinity):

what do you see know

OpenStudy (xapproachesinfinity):

forget about log thing that I said earlier just use exponent properties

OpenStudy (xapproachesinfinity):

\(\normalsize\color{blue}{ 4^{2x-3} =2^{2(2x-3)}=2^{4x-6} }\)

OpenStudy (xapproachesinfinity):

can you do the rest?

OpenStudy (xapproachesinfinity):

@jeshana

OpenStudy (anonymous):

Yes thank u

OpenStudy (xapproachesinfinity):

You're welcome. can you post your answer later

OpenStudy (anonymous):

I got 11/4 = 2.75 @xapproachesinfinity

OpenStudy (xapproachesinfinity):

Correct! Good job!